# Effects of interlayer Dzyaloshinskii–Moriya interaction on the shape and dynamics of magnetic twin-skyrmions

### v1

@@ Abstract
Skyrmions in synthetic antiferromagnets, constructed from ferromagnetic layers coupled by the interlayer exchange interaction (IEC), provide an example of such 3D structures  [11, 12, 13, 14, 15, 16]. A two-dimensional skyrmion is formed inside each layer, with the spin directions in the skyrmions being reversed between the layers due to the antiferromagnetic coupling. Skyrmions may often be characterized by two independent parameters [17]: the vorticity and the helicity. The vorticity is the integer winding number determining how many times and in which direction the in-plane spins wind around the out-of-plane spin in the center, which determines the topological charge $Q$ when also considering the direction of the center spin. The helicity angle characterizes the rotational sense of the spins when moving along the radial direction out from the center, including Bloch-type and Néel-type rotations. The total topological charge of synthetic antiferromagnetic skyrmions is zero due to the symmetric but opposite magnetization between the layers, which forces the skyrmions to move along the driving spin-polarized current in the current-in-plane (CIP) geometry, leading to a suppression of the skyrmion Hall effect [18]. The cancellation of the topological charge also strongly enhances the diffusion in these systems [19]. The helicity influences the direction of skyrmion motion in the current-perpendicular-to-plane (CPP) geometry as well [20, 21, 22], which includes the effect of the so-called spin–orbit torque [23]. Quantum helicity eigenstates have been proposed as a basis for skyrmion qubits [24]. The helicity has been demonstrated to vary in thick magnetic multilayers due to demagnetization effects [10], and it may also be influenced by strain and an external magnetic field [25].
The antisymmetric counterpart of the IEC is the interlayer Dzyaloshinskii–Moriya interaction (IL-DMI) [26, 27]. This chiral coupling leads to a non-zero chirality between the layers [28, 29, 30, 31, 32], which could facilitate field-free spin-orbit torque switching of the magnetization [33]. However, the influence of the IL-DMI on the equilibrium structure and the dynamics of skyrmions in multilayers remains unexplored.
@@
Figure 1: Effect of IL-DMI on Néel-type twin-skyrmions. A schematic of the physical system studied here is shown on the left, consisting of two magnetic layers and a non-magnetic layer between them. The contributions to the Hamiltonian in Eq. (7) from the different layers are written inside them. Without IL-DMI the material can host typical Néel-type skyrmions, as seen in (a). If the IL-DMI lies in the $xy$ plane, the twin-skyrmion becomes elongated along the direction of the IL-DMI vector, as seen in (b) and (c). Finally, if the IL-DMI points into the out-of-plane $z$ direction, the skyrmions in the two layers twist relative to each other by changing their helicities. The parameters are $J^{\text{IF}}=$12\text{\,}\mathrm{m}\mathrm{e}\mathrm{V}$\$, $|\mbox{$\mathrm{D}$}^{\text{IF}}|=$3\text{\,}\mathrm{m}\mathrm{e}\mathrm{V}$\$, $|\mbox{$\mathrm{J}$}^{\text{IL}}|=$0.6\text{\,}\mathrm{m}\mathrm{e}\mathrm{V}$\$, $B^{z}=$-3.5\text{\,}\mathrm{T}$\$. The IL-DMI is $D^{\textrm{IL}}=\mbox{$\mathrm{0}$}$ for (a), $D^{\textrm{IL}}=$8\text{\,}\mathrm{m}\mathrm{e}\mathrm{V}$\cdot\hat{\mbox{$\mathrm{x}$}}$ for (b), $D^{\textrm{IL}}=$8\text{\,}\mathrm{m}\mathrm{e}\mathrm{V}$\cdot\hat{\mbox{$\mathrm{y}$}}$ for (c), and $D^{\textrm{IL}}=$1\text{\,}\mathrm{m}\mathrm{e}\mathrm{V}$\cdot\hat{\mbox{$\mathrm{z}$}}$ for (d).
@@ Results
The possible effects of the IL-DMI on the shapes of the skyrmions are summarized in Fig. 1. Considering only the ferromagnetic IEC, the horizontal positions of the skyrmions in the two layers become locked to each other, but their shape remains unaffected, as shown in Fig. 1(a). Including the IL-DMI, a finite angle is opened between the spins above each other in the two layers. For two ferromagnetic layers, this angle is given by $\alpha=\arctan\left(|\mbox{$\mathrm{D}$}^{\text{IL}}|/J^{\text{IL}}\right)$ [35], and the magnetic moments will align in the plane perpendicular to the IL-DMI. Here, the skyrmions in the two layers deform differently to maximize their energy gain from the IL-DMI, and form an object which we call twin-skyrmion. For example, in Fig. 1(b) the IL-DMI points in the $x$ direction, preferring a tilting between the two layers in the $yz$ plane. The out-of-plane polarized background in the top layer tilts towards the positive $y$ direction, while in the bottom layer it tilts towards the negative $y$ direction, as shown by the white arrows in Fig. 1(b). For a Néel-type skyrmion shown in the figure, the spins above and below the skyrmion center point along the positive or negative $y$ direction, and the system gains energy by enlarging these areas, resulting in an elongation along the $x$ direction and a small shift of the skyrmion centers in the two layers oppositely along the $y$ direction. When the IL-DMI points along the $y$ direction instead, the polarized background tilts towards the negative and positive $x$ directions in the first and second layers, and the twin-skyrmion becomes elongated along the $y$ direction, as shown in Fig. 1(c). For Bloch-type skyrmions, the in-plane spin components are rotated by $90$ degrees; consequently, the elongation can also be observed perpendicular to the direction of the IL-DMI. The direction of elongation may be generally expressed as
@@
Figure 3: Change of twin-skyrmion size and helicity for out-of-plane IL-DMI. (a) Difference in helicity between the layers $\psi_{1}-\psi_{2}$ as a function of $D^{\textrm{IL}}_{z}$ and $J^{\textrm{IL}}$. Panel (b) shows line cuts along the lines colored correspondingly in panel (a); crosses represent simulation results while the line shows the numerical solution of the equations in the continuum limit, see Supplementary Note 1. (c) Twin-skyrmion size for the same parameters, with line cuts shown in panel (d). The system parameters are $J^{\text{IF}}=$6\text{\,}\mathrm{m}\mathrm{e}\mathrm{V}$\$, $|\mbox{$\mathrm{D}$}^{\text{IF}}|=$1.5\text{\,}\mathrm{m}\mathrm{e}\mathrm{V}$\$, $D_{x}^{\text{IL}}=D_{y}^{\text{IL}}=0$, and $B^{z}=$3.0\text{\,}\mathrm{T}$\$. For relaxing the structures, the damping parameter $\alpha=1$ was used, and the simulation was run for $5\times 10^{6}$ time steps, i.e., $10\text{\,}\mathrm{n}\mathrm{s}$.
@@
We investigate how the velocity of the twin-skyrmion is influenced by the direction and the strength of the IL-DMI. The velocity is characterized by its magnitude and direction, the latter usually expressed in terms of the skyrmion Hall angle, which is typically defined as the angle between the current driving the skyrmion and the velocity. In our simulations, we fix the polarization vector $\mathrm{P}$ along the $x$ direction, which in the spin–orbit torque picture corresponds to an electric current $\mbox{$\mathrm{j}$}_{\textrm{el}}$ flowing along the positive $y$ direction. We have observed above that the IL-DMI changes the shape and size of the skyrmions, and its out-of-plane component induces a difference in the helicity between the layers. The deformation and helicity of the twin-skyrmion has no effect on the topological charge $Q$ and therefore on the gyrocoupling vector $\mathrm{G}$. The dissipation tensor $\mathfrak{D}$ depends on the size and shape of the skyrmions, but is independent of the helicity. The tensor $\mathcal{B}$ depends on all of the mentioned parameters. For circularly symmetric skyrmions, the latter may be expressed as a function of helicity $\psi$ and vorticity $m$ as [21],
@@
The influence of the out-of-plane component of the IL-DMI is investigated in Fig. 5. As it was shown in Fig. 3, this type of IL-DMI causes a difference in the helicity between the skyrmions in the two layers. This would induce different velocity directions for the two parts of the twin-skyrmion, which is expected to cause an instability. This can also be observed in the simulations, because only for small IL-DMI values ($\mbox{$\mathrm{D}$}^{\text{IL}}_{z}<$0.6\text{\,}\mathrm{m}\mathrm{e}\mathrm{V}$\$ for the chosen parameters) is the motion of twin-skyrmions stable. It could be expected that the velocity of the twin-skyrmion is determined by the average of the velocity vectors in the two layers, thus the different velocity directions between the layers leads to a reduction of the net velocity. On the contrary, an increase in the velocity with $D_{z}^{\textrm{IL}}$ may be observed in Fig. 5. This is caused by an increase in the twin-skyrmion size with IL-DMI, as shown in Figs. 3 and 5. The force calculated from the $\mathcal{B}$ tensor increases approximately linearly with skyrmion size; see Supplementary Note 3 for the derivation. Since the changes in the dissipation tensor $\mathfrak{D}$ are small compared to $\mathcal{B}$, the velocity of the twin-skyrmion correlates very closely with the skyrmion size and the force acting on the skyrmion. Since the dissipation tensor depends weakly on the IL-DMI in this case, the Hall angle only increases by 0.3% when increasing $D^{\text{IL}}_{z}$ from 0.0 $0\text{\,}\mathrm{m}\mathrm{e}\mathrm{V}$ to $0.6\text{\,}\mathrm{m}\mathrm{e}\mathrm{V}$.
Figure 5: Effect of the $D^{\text{IL}}_{z}$ on the twin-skyrmion velocity and size. Values of $D^{\text{IL}}_{z}$ greater than $0.6\text{\,}\mathrm{m}\mathrm{e}\mathrm{V}$ do not result in stable motion. The system parameters are: $J^{\text{IL}}=$1.75\text{\,}\mathrm{m}\mathrm{e}\mathrm{V}$\$, $D^{\text{IL}}_{x}=D^{\text{IL}}_{y}=0$, $\mbox{$\mathrm{B}$}=$1.4\text{\,}\mathrm{T}$\$, $\alpha=0.1$, and $\beta_{d}=$1\text{\,}\mathrm{n}\mathrm{s}^{-1}$\$.

### v2

@@ Introduction
Skyrmions in synthetic antiferromagnets, constructed from ferromagnetic layers coupled by the interlayer exchange coupling (IEC), provide an example of such 3D structures  [10, 50, 27, 41, 21, 7]. A two-dimensional skyrmion is formed inside each layer, with the spin directions in the skyrmions being reversed between the layers due to the antiferromagnetic coupling. The effect of ferromagnetic IEC on skyrmion size and stability has been studied in Refs. [8, 43]. Skyrmions may often be characterized by two independent parameters [38]: the vorticity and the helicity. The vorticity is the integer winding number determining how many times and in which direction the in-plane spins wind around the out-of-plane spin in the center, which determines the topological charge $Q$ when also considering the direction of the center spin. The helicity angle characterizes the rotational sense of the spins when moving along the radial direction out from the center, including Bloch-type and Néel-type rotations. The total topological charge of synthetic antiferromagnetic skyrmions is zero due to the symmetric but opposite magnetization between the layers, which forces the skyrmions to move along the driving spin-polarized current in the current-in-plane (CIP) geometry, leading to a suppression of the skyrmion Hall effect [52]. The cancellation of the topological charge also strongly enhances the diffusion in these systems [11]. The helicity influences the direction of skyrmion motion in the current-perpendicular-to-plane (CPP) geometry as well [19, 48, 37], which includes the effect of the so-called spin–orbit torque [32]. Quantum helicity eigenstates have been proposed as a basis for skyrmion qubits [40]. The helicity has been demonstrated to vary in thick magnetic multilayers due to demagnetization effects [26], and it may also be influenced by strain and an external magnetic field [30].
The antisymmetric counterpart of the IEC is the interlayer Dzyaloshinskii–Moriya interaction (IL-DMI) [12, 36], which gathered recently a lot of attention in the research community [53]. This chiral coupling leads to a non-zero chirality between the layers [46, 18, 17, 2, 6], which could facilitate field-free spin-orbit torque switching of the magnetization [28]. However, the influence of the IL-DMI on the equilibrium structure and the dynamics of skyrmions in multilayers remains unexplored.
@@
Figure 1: Effect of IL-DMI on Néel-type twin-skyrmions. A schematic of the physical system studied here is shown on the left, consisting of two magnetic layers and a non-magnetic layer between them. The contributions to the Hamiltonian in Eq. (1) from the different layers are written inside them. Without IL-DMI the material can host typical Néel-type skyrmions, as seen in (a). If the IL-DMI lies in the $xy$ plane, the twin-skyrmion becomes elongated along the direction of the IL-DMI vector, as seen in (b) and (c). Finally, if the IL-DMI points into the out-of-plane $z$ direction, the skyrmions in the two layers twist relative to each other by changing their helicities. The parameters are $J^{\text{IF}}=$12\text{\,}\mathrm{m}\mathrm{e}\mathrm{V}$\$, $|\mbox{$\mathrm{D}$}^{\text{IF}}|=$3\text{\,}\mathrm{m}\mathrm{e}\mathrm{V}$\$, $|\mbox{$\mathrm{J}$}^{\text{IL}}|=$0.6\text{\,}\mathrm{m}\mathrm{e}\mathrm{V}$\$, $B^{z}=$-3.5\text{\,}\mathrm{T}$\$. The IL-DMI is $\mbox{$\mathrm{D}$}^{\textrm{IL}}=\mbox{$\mathrm{0}$}$ for (a), $\mbox{$\mathrm{D}$}^{\textrm{IL}}=$8\text{\,}\mathrm{m}\mathrm{e}\mathrm{V}$\cdot\hat{\mbox{$\mathrm{x}$}}$ for (b), $\mbox{$\mathrm{D}$}^{\textrm{IL}}=$8\text{\,}\mathrm{m}\mathrm{e}\mathrm{V}$\cdot\hat{\mbox{$\mathrm{y}$}}$ for (c), and $\mbox{$\mathrm{D}$}^{\textrm{IL}}=$1\text{\,}\mathrm{m}\mathrm{e}\mathrm{V}$\cdot\hat{\mbox{$\mathrm{z}$}}$ for (d).
@@ Results
Considered system
We describe the magnetic system by a classical atomistic spin model of two layers of a square lattice containing $128\times 128$ sites on top of each other. We split the spin Hamiltonian of our system into three parts,
$H=H_{1}+H_{2}+H_{\text{inter}},$ (1)
$-\mu_{s}\sum_{i}\mbox{$\mathrm{B}$}\cdot\mbox{$\mathrm{S}$}_{i}^{l},$ (2)
Inside each layer, we consider the nearest-neighbor interfacial exchange interaction $J^{\text{IF}}$; the nearest-neighbor interfacial DMI (IF-DMI) vector $\mbox{$\mathrm{D}$}^{\text{IF}}_{ij}$ perpendicular to the bonds between nearest neighbors preferring a Néel-type rotation, i.e., $\mbox{$\mathrm{D}$}^{\text{IF}}_{ij}=D^{\text{IF}}\mbox{$\mathrm{\hat{e}}$}_{z}\times\mbox{$\mathrm{\hat{r}}$}_{ij}$, where $\mbox{$\mathrm{\hat{r}}$}_{ij}$ is the normalized vector connecting sites $i$ and $j$; and the Zeeman interaction with an external magnetic field $\mathrm{B}$. $\mu_{s}$ is the magnetic moment at a site and $\langle i,j\rangle$ is the sum over all sites $i$ and their nearest neighbors $j$. We include two interlayer couplings mediated by a non-magnetic spacer layer separating the two magnetic layers: the IEC $J^{\text{IL}}$ and the IL-DMI $\mbox{$\mathrm{D}$}^{\text{IL}}$. If not specified otherwise, we consider the parameters $J^{\text{IF}}=$6\text{\,}\mathrm{m}\mathrm{e}\mathrm{V}$\$, $|\mbox{$\mathrm{D}$}^{\text{IF}}|=$1.5\text{\,}\mathrm{m}\mathrm{e}\mathrm{V}$\$, and $\mu_{s}=3\mu_{B}$, where $\mu_{B}$ is the Bohr magneton. These values lead to the stabilization of Néel-type isolated metastable skyrmions in a collinear background inside a single layer for a sufficiently high external magnetic field applied along the out-of-plane $z$ direction, as illustrated in Fig. 1(a). The value of $J^{\text{IL}}$ is typically varied between $1$ and $3$ meV, while $D^{\text{IL}}$ usually takes values between 1-1 and $1$ meV. Similar ranges for the DMI parameters have been determined in Ref. [35] for Co/TM/Co(0001) heterostructures based on first-principles calculations, where TM is a single atomic layer of a transition metal Cu, Ag, Au or Pt. The interfacial exchange and DMI interactions have been demonstrated to be strongly affected by nonmagnetic capping layers, which are not explicitly included in the spin model; see, e.g., Ref. [44] for first-principles calculations on Co/Pt(111) covered by various 5dd elements of a single monolayer thickness, where similar $J^{\text{IF}}$ and $D^{\text{IF}}$ values have been reported to the ones used here.
To study the dynamics of the structures, we consider the current-perpendicular-to-the-plane (CPP) geometry. This describes the case when the multilayer is placed on a non-magnetic substrate with strong spin–orbit coupling. Besides the substrate being responsible for the interfacial DMI, an electric current flowing inside the substrate gives rise to a spin current flowing perpendicular to the layers, which exerts a spin–orbit torque on the magnetization [32]. The dynamics is described by the Landau-Lifshitz-Gilbert (LLG) equation [24, 14] including these torque terms [1, 52],
where $\alpha$ is the Gilbert damping and $\gamma$ is the absolute value of the gyromagnetic ratio. The effective field $\mbox{$\mathrm{B}$}^{\text{eff},l}_{i}$ is given by the derivative of the Hamiltonian with respect to $\mbox{$\mathrm{S}$}^{l}_{i}$, $\mbox{$\mathrm{B}$}^{\text{eff},l}_{i}=-\mu_{s}^{-1}\dfrac{\partial H}{\partial\mbox{$\mathrm{S}$}^{l}_{i}}$; see Methods for its explicit expression. $\mathrm{P}$ is the direction of the current polarization, which will be assumed to be perpendicular to both the electric current or electric field direction inside the nonmagnetic substrate and the flow direction of the spin current perpendicular to the surface [32], $\mbox{$\mathrm{P}$}\parallel\mbox{$\mathrm{j}$}_{\textrm{el}}\crossproduct\hat{\mbox{$\mathrm{z}$}}$. The effect of the current is split into a field-like torque proportional to $\beta_{s}$ and a damping-like torque proportional to $\beta_{d}$. Since the field-like torque acts exactly like an external magnetic field, it distorts the skyrmions but does not set them into motion [48]. Therefore, we simplify the analysis by setting $\beta_{f}=0$ and only investigating the effect of the damping-like term driving the skyrmions.
@@
The possible effects of the IL-DMI on the shapes of the skyrmions are summarized in Fig. 1. The direction of the IL-DMI vector is highly dependent on the symmetry of the system. Both in-plane [9, 35, 18] and out-of-plane [2, 39] directions have been reported, depending on the symmetry breaking introduced in the heterostructure. Here, we investigate the effect of both components. Considering only the ferromagnetic IEC, the horizontal positions of the skyrmions in the two layers become locked to each other, but their shape remains unaffected, as shown in Fig. 1(a). Including the IL-DMI, a finite angle is opened between the spins above each other in the two layers. For two ferromagnetic layers, this angle is given by $\alpha=\arctan\left(|\mbox{$\mathrm{D}$}^{\text{IL}}|/J^{\text{IL}}\right)$ [35], and the magnetic moments will align in the plane perpendicular to the IL-DMI. Here, the skyrmions in the two layers deform differently to maximize their energy gain from the IL-DMI, and form an object which we call twin-skyrmion. For example, in Fig. 1(b) the IL-DMI points in the $x$ direction, preferring a tilting between the two layers in the $yz$ plane. The background polarized along the negative $z$ direction in the top layer tilts towards the positive $y$ direction, while in the bottom layer it tilts towards the negative $y$ direction, as shown by the white arrows in Fig. 1(b). For a Néel-type skyrmion shown in the figure, the spins above and below the skyrmion center point along the positive or negative $y$ direction, and the system gains energy by enlarging these areas, resulting in an elongation along the $x$ direction and a small shift of the skyrmion centers in the two layers oppositely along the $y$ direction. Note that a high value of the IL-DMI $D^{\textrm{IL}}=$8\text{\,}\mathrm{m}\mathrm{e}\mathrm{V}$\$ was chosen in this figure to highlight the elongation. When the IL-DMI points along the $y$ direction instead, the polarized background tilts towards the negative and positive $x$ directions in the first and second layers, and the twin-skyrmion becomes elongated along the $y$ direction, as shown in Fig. 1(c). For Bloch-type skyrmions, the in-plane spin components are rotated by $90$ degrees; consequently, the elongation can also be observed perpendicular to the direction of the IL-DMI. The direction of elongation may be generally expressed as
@@
Figure 3: Change of twin-skyrmion size and helicity for out-of-plane IL-DMI. (a) Difference in helicity between the layers $\psi_{1}-\psi_{2}$ as a function of $D^{\textrm{IL}}_{z}$ and $J^{\textrm{IL}}$. Panel (b) shows line cuts along the lines colored correspondingly in panel (a); crosses represent simulation results while the line shows the numerical solution of the equations in the continuum limit, see Supplementary Note 1 and Supplementary Fig. S1. (c) Twin-skyrmion size for the same parameters, with line cuts shown in panel (d). The system parameters are $J^{\text{IF}}=$6\text{\,}\mathrm{m}\mathrm{e}\mathrm{V}$\$, $|\mbox{$\mathrm{D}$}^{\text{IF}}|=$1.5\text{\,}\mathrm{m}\mathrm{e}\mathrm{V}$\$, $D_{x}^{\text{IL}}=D_{y}^{\text{IL}}=0$, and $B^{z}=$3.0\text{\,}\mathrm{T}$\$. For relaxing the structures, the damping parameter $\alpha=1$ was used, and the simulation was run for $5\times 10^{6}$ time steps, i.e., $10\text{\,}\mathrm{n}\mathrm{s}$.
@@
We investigate how the velocity of the twin-skyrmion is influenced by the direction and the strength of the IL-DMI. The velocity is characterized by its magnitude and direction, the latter usually expressed in terms of the skyrmion Hall angle, which is typically defined as the angle between the current driving the skyrmion and the velocity; see Supplementary Note 2. In our simulations, we fix the polarization vector $\mathrm{P}$ along the $x$ direction, which in the spin–orbit torque picture corresponds to an electric current $\mbox{$\mathrm{j}$}_{\textrm{el}}$ flowing along the positive $y$ direction. We have observed above that the IL-DMI changes the shape and size of the skyrmions, and its out-of-plane component induces a difference in the helicity between the layers. The deformation and helicity of the twin-skyrmion has no effect on the topological charge $Q$ and therefore on the gyrocoupling vector $\mathrm{G}$. The dissipation tensor $\mathfrak{D}$ depends on the size and shape of the skyrmions, but is independent of the helicity. The tensor $\mathcal{B}$ depends on all of the mentioned parameters. For circularly symmetric skyrmions, the latter may be expressed as a function of helicity $\psi$ and vorticity $m$ as [48],
@@
The influence of the out-of-plane component of the IL-DMI is investigated in Fig. 5. As it was shown in Fig. 3, this type of IL-DMI causes a difference in the helicity between the skyrmions in the two layers. This would induce different velocity directions for the two parts of the twin-skyrmion, which is expected to cause an instability. This can also be observed in the simulations, because only for small IL-DMI values ($\mbox{$\mathrm{D}$}^{\text{IL}}_{z}<$0.6\text{\,}\mathrm{m}\mathrm{e}\mathrm{V}$\$ for the chosen parameters) is the motion of twin-skyrmions stable. It could be expected that the velocity of the twin-skyrmion is determined by the average of the velocity vectors in the two layers, thus the different velocity directions between the layers leads to a reduction of the net velocity. On the contrary, an increase in the velocity with $D_{z}^{\textrm{IL}}$ may be observed in Fig. 5. This is caused by an increase in the twin-skyrmion size with IL-DMI, as shown in Figs. 3 and 5. The force calculated from the $\mathcal{B}$ tensor increases approximately linearly with skyrmion size; see Supplementary Note 3 for the derivation and Supplementary Fig. S2. Since the changes in the dissipation tensor $\mathfrak{D}$ are small compared to $\mathcal{B}$, the velocity of the twin-skyrmion correlates very closely with the skyrmion size and the force acting on the skyrmion. Since the dissipation tensor depends weakly on the IL-DMI in this case, the Hall angle only increases by 0.3% when increasing $D^{\text{IL}}_{z}$ from 0.0 $0\text{\,}\mathrm{m}\mathrm{e}\mathrm{V}$ to $0.6\text{\,}\mathrm{m}\mathrm{e}\mathrm{V}$.
Figure 5: Effect of the $D^{\text{IL}}_{z}$ on the twin-skyrmion velocity and size. Values of $D^{\text{IL}}_{z}$ greater than $0.6\text{\,}\mathrm{m}\mathrm{e}\mathrm{V}$ do not result in stable motion. The system parameters are: $J^{\text{IL}}=$1.75\text{\,}\mathrm{m}\mathrm{e}\mathrm{V}$\$, $D^{\text{IL}}_{x}=D^{\text{IL}}_{y}=0$, $\mbox{$\mathrm{B}$}=$1.4\text{\,}\mathrm{T}$\$, $\alpha=0.1$, and $\beta_{d}=$1\text{\,}\mathrm{n}\mathrm{s}^{-1}$\$. The simulation ran for $2\times 10^{6}$ time steps, i.e., $2\text{\,}\mathrm{n}\mathrm{s}$.